### v2

# Laser-FLASH: radiobiology at high dose, ultra-high dose-rate, single pulse laser-driven proton source

Abstract
Laser-driven proton sources have long been developed with an eye on their
potential for medical application to radiation therapy. These sources are compact,
versatile, and show peculiar characteristics such as extreme instantaneous dose
rates, short duration and broad energy spectrum. Typical temporal modality of
laser-driven irradiation, the so-called fast-fractionation, results from
the composition of multiple, temporally separated, ultra-short dose fractions.
In this paper we present the use of a high-energy laser system for delivering the
target dose in a single nanosecond pulse, for ultra-fast irradiation of
biological samples. A transport line composed by two permanent-magnet quadrupoles
and a scattering system is used to improve the dose profile and to control the
delivered dose-per-pulse. A single-shot dosimetry protocol for the broad-spectrum
proton source using Monte Carlo simulations was developed. Doses as high as
could be delivered in a single shot, lasting less than
over a diameter sample holder, at a dose-rate exceeding
. Exploratory application of extreme laser-driven irradiation
conditions, falling within the FLASH irradiation
protocol[1], are presented for in vitro and in vivo irradiation. A reduction of radiation-induced oxidative
stress in-vitro and radiation-induced developmental damage in vivo were observed,
whereas anti-tumoral efficacy was confirmed by cell survival assay.
Introduction
Radiation therapy is a cornerstone in cancer management. Apart from X-rays, which
represent the strong majority of treatments, other radiation qualities and disparate
spatial or temporal source parameters are used to match particular therapeutic needs.
Unlike X-rays, protons have a rapid distal dose fall-off and a reduced proximal dose.
Laser-driven proton sources have been proposed as a promising alternative to conventional
(cyclotrons, synchro-cyclotrons)
accelerators[2, 3].
Lasers, as an energy source, do not require specific radioprotection; moreover ion plasma
sources produce broad spectra, which could offer novel alternative strategies for
producing a spread-out Bragg peak (SOBP), used for the treatment of deep-seated tumors.
As of today, existing laser technologies and explored acceleration strategies do not
provide sufficient kinetic energies for the medical application, although successful
experiment of in-vitro[4, 5] and surface
in-vivo[6] irradiation made a huge step forward
to demonstrate their practical use in radiation biology.
The recent discovery of the FLASH effect by Favaudon et al.[7] renewed the interest around the role of dose-rate and temporal modality of dose deposition in defining the biological and the physiological effect of ionizing radiation. Some advantages demonstrated for in-vivo and human irradiation (see Wilson et al.[8] and references therein for a consistent review) suggest a possible improvement of the therapeutic window at high dose rate, which motivates the exploration of non-conventional temporal dose deposition modalities. Laser-driven particle sources produce short and bright particle bunches, as each laser pulse independently extracts and accelerates a short and bright packet. Owing to the ultra-fast nature of laser pulses at relativistic intensity, laser-accelerated protons have a duration, at the source, between few picoseconds and few nanoseconds[9, 10].
An ever increasing number of laboratories worldwide are exploring the biological effects of laser-accelerated protons, where relevant irradiation conditions are met in terms of particle penetration, irradiated volume and average dose-rate. Such systems are often limited to an energy-per-pulse of a few to tens of joules and repetition rates in the Hz range or lower. The total useful proton charge from such systems is sufficient to deposit a dose of a few to a fraction of per shot, where a projected target surface of is considered. In this condition, multiple laser pulses are required to produce a target dose within to , with a total irradiation time not shorter than a few seconds and up to a few minutes. This irradiation modality, where the dose is deposited by separate fractions at ultra-high dose-rate, termed “fast fractionation”, was shown to produce particular biological effects in some conditions[4, 11].
In this paper we present the use of a kJ-class laser to achieve the total target dose
within a single laser-driven proton pulse. A set of permanent-magnet quadrupoles is used
to transport and shape the protons to an irradiation area in air, far from the
interaction point. This technique gives access to doses as high as
delivered within a few nanoseconds over a in-vitro sample.
The transport line provides an additional control on the irradiation conditions, such as the
ability to vary the projected charge density in the proton bunch, for performing
dose-escalation experiments. Moreover, the spectral selection operated by the quadrupoles
can be used to produce a uniform depth dose deposition (SOBP) within thicker targets
(up to of water).
In order to correctly apply our particle source to radiobiology experiments, a dosimetry
protocol for the reconstruction of the spectral content of each laser shot is developed
and applied to all of the explored irradiation conditions.
A survival assay of human glyoblastoma cell line U87-MG is performed as a confirmation of the
toxicity to cancerous cells of the irradiation protocol. The generation of stress-dependent oxydative stress is studied
in-vitro in healthy (MRC5) and tumoral (U87-MG) cell lines, showing reduced DNA damage
in the healthy cells. This suggests a protective effect of the high-dose-rate irradiation
condition. Evaluation of zebrafish embryos development following irradiation under these
conditions suggest onset of the FLASH effect with short, laser-driven, single
pulse proton irradiation scheme.
Source and transport
The experiment was performed in two separate campaigns at the LULI laboratory (École Polytechnique, Palaiseau, France) on the pico2000 laser system. The picosecond beam (duration , energy per pulse, ) is focused by a off-axis parabola (PFL , d=).
Protons are accelerated in the target-normal sheath acceleration (TNSA) scheme. Optimal acceleration conditions, showing both highest charge and cut-off energy, were determined from previous experimental campaigns[12]. The TNSA mechanism relies on the energy transfer from the laser-heated electron fraction to the cold ion plasma, during the plasma expansion in vacuum (see Macchi 2017[13] and references therein); ion spectrum is thermal with a high-energy cutoff. In order to characterize the particle source for magnetic transport design, a set of stacked radiochromic films (RCF) is used. The stack is installed at from the interaction point () and irradiated with protons from a single laser shot. The spectral and spatial features of the accelerated beam are extracted from the irradiated RCFs following an exponential fitting routine (described in section 3.1.2 of Cavallone, 2020[14]) which solves with higher precision the cutoff energy and the overall spectral shape. The analysis procedure also enables a rough estimate of the total available charge at the source.
A calibration shot to assess the source charge and spectral features is illustrated in Figure 2a for a thick gold target. The spectrum exhibits the expected exponential profile, with a cutoff at . Radiochromic films # through # were saturated. Different spectral components exhibit a decreasing divergence at increasing energy (Figure 2b). It’s interesting to note that protons with energy are emitted with constant divergence, producing a rather sharp boundary on the radiochromic films deeper in the stack. At higher energies, the beam divergence decreases following a parabolic profile. According to the general calibration function of the HD810 film, the total charge within the exponential fit for the presented shot exceeds .
Transport configuration
The transport line is designed to produce a controlled irradiation to a plane outside the
experimental chamber. Air-vacuum separation is guaranteed by a thick Kapton
film, situated at from the interaction point, whereas the irradiation
plane lays at (Figure 1). The transport line is composed
by two permanent-magnet quadrupoles and a variable-thickness scattering filter.
Quadrupoles Q1 and Q2 are respectively and
long, and have field gradients of and
. Quadrupoles construction and calibration is detailed in
Rösch et al.[15, 16]. Both
quadrupoles have an inner bore of diameter; considering the shortest possible
distance between quadrupole Q1 and the proton source,
, the bore sets a maximum angular acceptance of .
The scattering system (ScS) is composed by four different selectable aluminum
filters of varying thickness to be
inserted between the quadrupole Q2 and the Kapton window, to provide diffusion of the proton beam,
improving divergence and transverse uniformity.
All elements on the transport line are motorized and equipped with optical encoders,
enabling reliable positioning even in presence of strong magnetic forces between the
elements.
The role of the transport line is to control the particle density and, consequently, the deposited dose at the target plane, far from the laser-plasma interaction point. During the design step the deposited dose at the irradiation plane is simulated by Monte Carlo methods (Geant4 toolkit) for different beam line configurations. The source is modeled as being purely thermal with a high energy cutoff and no angular dependence. This assumption is based on the small effective aperture of the first quadrupole, which acts as a filter on the broad divergent source. The largest accepted half-width angle () is, according to measurements in Figure 2, smaller than the source divergence at the cutoff energy. For this reason we decided to drop the angular dependence in the source modelling; the useful spectral charge is then represented analytically as:
where E is the proton kinetic energy, E0 the spectral distribution temperature and Q a charge. The term Q (measured
in nC∗ from now on to avoid confusion) indicates that the integrated charge of the
source term expressed in equation (1) differs from the total charge at the
laser-plasma source, as it is limited to the charge emitted within the first quadrupole
input acceptance angle (see Monte Carlo simulation section in Methods for
further details).
Two sets of configurations are designed for the particle transport. A first condition (termed dose escalation, DE) is studied for in-vitro irradiation. The surface to be irradiated is set to have a diameter of , centered on axis on a Lumox capsule and the biological sample represented by a thick water volume. In this condition the quadrupole Q2 is set at from the first quadrupole. In this configuration the scattering filter holds a aluminum foil and is moved to vary the particle density at the irradiation plane, hence the deposited dose within the projected surface (Figure 3).
An additional configuration was studied for in vivo irradiation (Zebrafish embryos, ZF). In this condition the target dose at the sample is set to be as close as possible to a well defined condition (, as in Bourhis et al.[17]) while keeping the maximum in-depth (SOBP) uniformity. This irradiation condition was obtained with a scattering filter of positioned at a distance of past the quadrupole Q2; in this configuration a smaller profile and a higher dose is produced at the irradiation plane, as depicted in Figure 4. The corresponding region-of-interest (ROI) is here limited to a diameter of . .
Single pulse irradiation and dosimetry
In our experimental condition, the target dose is deposited by the charge driven by a
single laser pulse, within a time shorter than . During irradiation it is not
possible to monitor or regulate the deposited dose with a monitor chamber, as performed
during previous
experiments[18].
Spatially resolved dose maps
are recorded by calibrated EBT3-XD radiochromic films as a reference to
the dose in the target sample. However, the width of the proton spectrum and the low
overall kinetic energy in the beam make the energy loss in the RCF itself non negligible.
Consequently, not only the dose at the RCF differs from the dose at the target
sample, but the sole presence of the RCF does modify the target irradiation conditions.
The shot-to-shot variation of the laser-driven proton source parameters (energy, charge)
we observed did not enable to establish one irradiation condition within an acceptable
error margin beforehand. Laser-driven proton acceleration in the TNSA scheme is in fact
very sensitive to the total energy and to the laser-temporal contrast, which are beyond
the user control.
In our case, irradiation driven by a single laser pulse, hinders the ability to define an
average spectrum for calibrating the response radiochromic films on the irradiation line.
TNSA Proton source reconstruction
Spectrum and charge changes at the source result in variations of the total deposited dose, as well as in the dose distribution among different planes in the irradiated volume (see for example the in-vitro holder structure Figure 5a). One fundamental assumption on the shot-to-shot behaviour of the proton source is the conservation of its exponential nature, which is well justified for the TNSA acceleration mechanism. Under this hypothesis, shot-to-shot parameter change can be considered to affect the only two free parameters in the spectral distribution (1), namely the total charge Q and the spectral temperature E0. A third parameter known to depend on laser conditions is the cutoff energy, Ehigh, which is usually defined as a reference parameter for laser-driven proton acceleration conditions. For our purposes, a simple rejection criterion can be set for those shots whose cutoff energy is not sufficient to penetrate the irradiated target stack. Minor variations can be neglected because, according to Monte Carlo modelling of the irradiation line, the highest portion of the spectrum contributes to a lesser extent to the total deposited dose, with respect to the central portion of the spectrum. For these reasons we consider the cutoff energy constant throughout our analysis.
At least two (although often three of them) RCF were used in most experimental events, inserted before and after the biological sample during irradiation (see panel (a) in Figure 5). Deposited dose in all of the RCFs and the irradiated target is simulated for the corresponding beam-line configuration and sources with a varying temperature E0 of the thermal spectrum between . Following previous considerations and according to equation (1), total charge Q and spectral temperature E0 can be extrapolated from the Monte Carlo simulation which would correctly predict the dose ratio between the superposed RCF.
Simulated deposited dose is scaled to the measured dose on radiochromic films for all the temperatures in the set. The corresponding charge, averaged for the available RCFs in a single shot, can be then expressed as:
where is the experimental integrated dose on the i-th film in the target and is the simulated dose at the spectral temperature E0 on the i-th film. Figure 5a,b show, as an example, the relationship between doses at different planes for the case of an experimental shot in the dose escalation configuration with filter at . From the analysis of the calculated dose ratios in the different simulations, it is possible to set the matching temperature by minimizing the standard deviation of the extrapolated charge. As an example Figure 7 depicts the minimization of the ratio in the case of a test shot, which points to a temperature of . The corresponding value of is then used to rescale the simulation and assess the dose in the biological sample. The obtained values are shown in the table 6.
| Position in the stack | RCF 1 | RCF 2 | Sample | RCF 3 |
| Dose () |
In some of the experimental conditions only one RCF foil was used in front of the
biological target. This was the case for fish irradiation, where we aimed at preserving the best
possible SOBP uniformity; adding a second radiochromic film would have reduced the penetration
depth of the proton beam. Furthermore the actual design of the ZF holder does not allow the insertion
of a back RCF film (see Figure 15 for details).
In these cases only one experimental point is available for assessing the ratio between
the deposited dose and the RCF reading, hence it is not possible to determine the
parameters E0 and Q in a unique way using the previously described protocol. In
order to assess the ratio between the RCF reading and the deposited dose we decided to use the temperature
distribution from the in-vitro events,
shown in Figure 8b as a statistical weight to the depth-dose curve
and ratio to RCF depicted in Figure 8a.
Following this method, the quantities sample-dose-to-RCF and depth-dose-error are calculated as
where the weights (spectrum temperature occurrence probability) are represented in Figure 8.
| Temperature E0 (MeV) | Dose (Gy/nC∗) | Depth-dose error (relative) | RCF Dose (Gy/nC∗) | Dose ratio |
Application to radiation biology
The irradiation beam-line was used for irradiation on several in-vitro and in-vivo
biological targets, in
order to verify its functional hypothesis, provide a rough validation of the dosimetry
protocol and start to explore this novel irradiation condition.
In the following, radiation qualities are indicated as SP-LDP for “single-pulse
laser-driven protons”, FF-LDP for “fast-fractionated laser-driven protons”
(indicating a dose deposited by multiple separated fraction at ultra-high instantaneous
dose-rate) and CAP for “conventional-accelerated protons”.
Cell survival assay after exposure to laser-driven protons (LDP) at pico2000
The impact of SP-LDP is evaluated on the highly resistant
glioblastoma cells of the line U87-MG, through a cell-survival assay. Cells are prepared
as in a previous experimental campaign[4] where the
effect of laser-pulse pacing in a fast-fractionation irradiation modality was explored. It’s
worth noting that U87-MG showed no sensitivity to the average dose-rate at fixed dose
within the range explored during previous experiments (dose-per-fraction:
, delay between fractions: ).
U87-MG cells were irradiated in single pulse modality with doses ranging from to
and the resulting dose-response survival curves obtained from
non-clonogenic survival assays (see Methods). Figure 10 show a superposition
between novel SP-LDP points and those already published in Bayart et al., 2019
[4]. The D10 values resulting from a
linear-quadratic model fit do show good
agreement between the three experiments, confirming the toxicity of our irradiation
conditions on cancerous cells and the validity of the dose escalation scheme. The set of
values across the two experiments is summarized in
Table 11, along with average and instantaneous dose-rate
(respectively and ) and the value
of the fit.
Oxidative stress-dependent DNA damage in healthy and tumoral cell lines after exposures to SP-LDP.
In the described conditions, the total dose is deposited in a time shorter than , at a dose-rate exceeding . These conditions are several orders of magnitude shorter and more intense than those indicated as a threshold for FLASH effect in Bourhis et al, 2019[17]. In this single-pulse condition, dose deposition happens in a temporal span faster than the homogeneous chemical step[19]. Although the associated mechanistics is not yet known, FLASH effect has been associated to the reduction of oxidative stress in healthy tissue.
To address the question whether SP-LDP could have a different effectiveness to generate oxidative stress-dependent DNA damage, we aimed to measure 8-Hydroxy-2’-deoxyguanosine (8-OHdG). 8-OHdG is one of the most widely studied oxidized metabolites and is considered a biomarker for oxidative damage of DNA[20, 21]. Human healthy fibroblasts MRC5 and U87-MG cells were exposed to a target dose of of SP-LDP ( and in MRC5 and U87-MG cells respectively according to post-irradiation dosimetry). The amount of 8-OHdG was determined in cells harvested one hour later and analyzed by Elisa assay (see Methods). Using this detection method, the amount of 8-OHdG detected is inversely related to absorbance. Figure 12 shows the ratio of the values obtained with MCR5 cell line out of those of U87-MG. At the basal level, MRC5 and U87-MG cells present similar amounts of 8-OHdG, indicated by a ratio close to one. However, following irradiation with SP-LDP, 8-OHdG ratio increase significantly (, Kolmogorov-Smirnov test). These result suggests that, following exposure to ultra-high dose-rate protons, the glioblastoma cell line model receives a higher fraction of DNA damage due to oxidative stress than he healthy cell model. This is the first time that a link between DNA damage, oxidative stress and the differential response between healthy and tumor cells is observed. It may be also be an indication of an in vitro sparing effect at ultra-high dose-rate, similar to FLASH irradiation protocols.
Zebrafish embryo development evaluation as preclinical in-vivo model for radiation research at pico2000
Radiation-induced toxicity reduction at constant dose and higher dose-rate (FLASH effect) have been exclusively observed with in vivo models (for review see Vozenin et
al, 2022[1]) including mice, cats, dogs,
pigs and zebrafish (Danio rerio). Aiming at a validation of in vivo
irradiation of laser-driven protons, zebrafish do represent a good
candidate.
Zebrafish (ZF) embryos have a small size (), demanding lower proton
energy and relatively small SOBP. Embryos correspond to functionally and morphologically organisms with the size of
organoids, well recognized as a good model for radiobiological evaluation of ionizing
radiation, particularly suitable for lesser penetrating
laser-accelerated
protons[22, 23, 24].
In order to irradiate ZF embryos, the transport line was setup to produce a uniform
irradiation field over diameter and depth in water. A specific
holder was designed to confine the embryos in close contact with the irradiation field
within the irradiation field surface while
preserving their survival conditions (see Methods).
The holder allows the irradiation of several embryos at the same time. Embryos were
irradiated at post-fertilization, when the embryo has a diameter of roughly
. Developed fish were fixated post-irradiation, and the fish length measured
as an indicator of SP-LDP toxicity (Figure 13).
Irradiated animals exhibited changes in morphology and in the length due to spine
curvature and developmental deterioration.
As expected, after irradiation with SB-LDP, zebrafish embryos showed a significant
decrease in total length, with a length corresponding to
of non-irradiated animals. As the total dose was deposited in a
single bunch, our irradiation conditions fall within the current definition of FLASH. It
is yet to verify if the much shorter duration and the much higher instantaneous dose rate
do have measurable effects in addition to those already demonstrated.
FLASH effect on zebrafish embryos was observed in similar conditions ( dose
fraction at post-fertilization, developmental evaluation at
post-irradiation) in Bourhis et al. 2019
paper[17], where a developmental
improvement is observed between FLASH and conventional irradiation conditions.
In our experiment, length decrease measured
for irradiated embryos was about of the non irradiated for a slightly
higher dose (). This observation suggests that SP-LDP condition
could have triggered FLASH effect in the irradiated ZF embryos.
Discussion
Due to their favorable ballistic properties, hadrons (protons and carbon ions)
represent a better alternative for the radiation therapy of solid tumors affecting
organ-at-risk. The Bragg peak dose deposition profile allows the dose to be better
concentrated within the target volume, reducing side effects to neighboring healthy
tissues. The development of laser plasma technology and its new paradigms for protons
acceleration (Wilks et al., 2001) raise new possibilities for studying the effects of dose
delivery modalities[3].
However, due to the used particle (electrons or protons) acceleration technology
(laser-driven sources), the total deposited dose is delivered as a sequence of multiple
ultra-short separate fractions at ultra-high instantaneous dose rates; the dose per pulse
(hence the number of bunches required for a given dose) and the effective repetition rate
(hence the total irradiation time) depend on technical choices or limitations at the
facility[25, 26, 27, 28, 11, 29, 2].
In the past we the radiobiological impact of the repetition rate of bunches and
we showed that, at constant dose, the temporal dose deposition modality of LDP
is a key parameter in determining cancer cells response[4].
Since the first proof of concept of irradiating cells with
LDP[28, 11]
we have been able to explore the biological impact of
laser-accelerated protons of highly resistant glioblastoma cells, for which proton
therapy is one of the main indications. Cell
survival of U87-MG glioblastoma cell line showed, as expected, similar effectiveness of
single-pulse-LDP (SP-LDP), compared to fast fraction-LDP and conventional beams, to induce
cell killing (Figure 10). Indeed, we showed that this cell line was unsensitive to the
temporal parameter of dose deposition resulting from the absence of functional PARP1
protein which is probably responsible of its high radioresistance.
We also investigated the ability of laser-accelerated protons to generate
DNA double-strand breaks (DSBs); it is confirmed that LDP and conventional beams have
similar effectiveness, although a number of non-significant divergences were reported
(mainly arising from differences in experimental procedures and endpoints).
Flash effect and related healthy tissue sparing have been reported
to be linked to reduced production of reactive oxygen species
(ROS)[30, 19].
In this study, the high charge generated by the high energy Pico2000 laser, using
the well-established TNSA acceleration technique, enabled irradiation conditions
compatible with the requirements of the FLASH protocol. Owing to the high energy/low
repetition rate configuration, the FLASH condition was reached within a single
laser pulse; proton bunches duration is shorter than ,
which implies a dose-rate exceeding .
The use of a quadrupole transport line, enabled to control the deposited dose within the
single proton pulse. In our setup, a moving scattering filter installed after the
second magnetic focusing element was used to adjust the diffusion center of proton
emission, thereby affecting the particle density at the irradiation plane.
This technique enables dose escalation, even with the
ultra-fast nature of laser-driven acceleration mechanism. Additionally, the two focusing
elements naturally influence the particle spectrum by limiting the low energy end, thus
enhancing the dose contribution from the central portion. As a result, this effect
produced a more uniform SOBP through the thick biological sample.
We demonstrated how the temperature of the TNSA proton emission could be used as a
parameter to fit the available diagnostics on the proton transport line. In fact the
otherwise non-optimal acceleration repeatability proved to loosely depend on temperature
and strongly on the amount of accelerated charge. The Analysis protocol we set up, not
only enabled a precise reconstruction of the dose within the target volume, but also to
extract beam parameters otherwise non accessible. This strategy will be of use for future
experiments where TNSA accelerated protons will be used for practical applications, most
importantly for laser-driven single-pulse FLASH irradiation condition.
Notwithstanding the challenges of retrieving irradiation conditions post-irradiation, our experiment
highlights a fundamental limitation of laser-driven single-pulse FLASH: it is not
possible to monitor, and therefore to control, the total deposited dose before the
irradiation is completed.
The experimental evidence we gathered under laser-driven FLASH conditions was used to explore whether induction of oxidative stress-dependent DNA damage could differ between healthy and tumor cells under these irradiation modalities. By comparing the ratios between the two cell types, we observed that oxidative stress-dependent DNA damage was equivalent in non-irradiated cells; however it was higher in U87-MG glioblastoma cells than in healthy MRC5 fibroblasts (Figure 12). This result suggests that applying SP-LDP may be able to generate a greater amount of oxidative stress-dependent DNA damage in tumor model than in a healthy one. This hypothesis aligns with recent data indicating that FLASH irradiation produces significantly more ROS than conventional irradiation, which can be eliminated much more efficiently in normal tissues during steady-state metabolism than in tumors.
Acknowledgments
AF acknowledges region Ile-de-France for the contract Sesame/IDRA, and ANR contract FemtoDose. AF and CG acknowledge CNRS contract MITI/BioRapide.
Author contributions statement
AF proposed the experiment in collaboration with MC and LR. TR, LT, JS and KP provided the quadrupole system. AF, LR and EB performed the experiment with support from MC, JdM, TR, AP. LdM and AP provided calibration of RCF. EB, ILJ and CG prepared and analyzed biological samples. AF and EB wrote the paper with contribution from LR, TR, CG, LdM, AP, KP and JS.
Methods
Dosimetry
Monte Carlo simulation
The proton beamline is simulated using the Monte Carlo toolkit Geant4, where quadrupoles
are represented
by measured field maps. The possibility of simulating magnetic optics with TNSA
sources and energies using Monte Carlo methods was studied in greater detail and validated in
Cavallone et al., 2019[31] by comparing it
with codes that account for space charge.
Monte Carlo simulations are run with a point proton source with an exponential spectrum as
in equation (1), where lower and higher cutoffs matching experimental
values. The lower cutoff is set to , which
is well below the minimum energy needed to reach the first radiochromic film in absence of
a scattering filter; while the maximum is set to , which corresponds to the
highest measured energy. Source has isotropic angular distribution up to ; source temperature E0 and effective charge Q are varied to
match dosimetry measurement.
The total proton charge on the test shot was estimated to be Q, obtained from the exponential fitting algorithm described in
[14]. The effective charge ratio (i.e. the total charge within
the input acceptance angle of the first quadrupole) is calculated by numerical
integration from fitted curves in Figure 2 and found to
be which is used as an order of
magnitude of the number of particles entering the transport line. This figure does
not take in account the efficiency of the particle transport, nor the actual
charge within the ROI on the biological sample. Given the numerical complexity of
the transport line, the number of primaries in the Monte Carlo simulation is
reduced to (roughly a factor of the
experimental charge) and then scaled back to . A test simulation in
the dose escalation configuration shows that, on the average,
are recorded per each pixel at the first RCF
(Figure 14).
EBT radiochromic films are represented in the simulation by
of composite material with correct relative atomic abundances and density.
For a setup where multiple RCF are present, separate simulations are run where each RCF
material is switched to water, in order to account for the equivalent deposited
dose-to-water. In fact, the RCF calibration is obtained as a water
equivalent (density ) while the actual average density of the
radiochromic foil is , where and are respectively the
polyester and the sensitive layer volume fractions in a radiochromic EBT3 type foil. This
correction is needed in order to correctly simulate the dose deposition of low kinetic energy
protons in the sensitive regions of dosimeters and biological sample. The observed
correction due to the change water/polyesther can be as high as .
Radiobiology procedures
Cell culture
The human glioblastoma cell line U87-MG was cultured in Dulbecco’s modified Eagle’s minimum medium with Glutamax (Thermo Fisher Scientific). Cells were grown as monolayers, supplemented with fetal calf serum (PAA) and penicillin and streptomycin (ThermoFisher Scientific) in plastic tissue culture disposable flasks (TPP) at in a humidified atmosphere of in air.
Cell survival assay
The cell containers used for irradiation were lumox® dish 35 (SARSTEDT) exhibiting a thick lumox® bottom face. Depending on beam transverse profile and position observed on radiochromic films (ref) a circular area was delimited on the internal face of the lumox® membrane, where cells were seeded. Cells were let grow overnight in of medium. To generate dose-response survival curves U87-MG cell lines were subjected to doses varying from with single bunch of laser driven protons. Survival curves resulting from irradiation with CAP or fast-fractionation LDP were retrieved from [4]. After exposure to ionizing radiations, cells were incubated for in standard conditions. Cells were harvested with Accutase (Merck), dispatched into different wells of -well plates (TPP) in of medium and grown for five generations corresponding to for U87-MG cell line. Cells were harvested with Accutase which was then inactivated using an equal volume of PBS (ThermoFisher Scientific) supplemented with fetal calf serum. The final volume was adjusted to with PBS and of each well were dispatched into a non-sterile U-bottom -well plate (TPP). In each well, of a propidium iodide solution (Sigma, 100 in 1X PBS) were added just before flow cytometry counting. Cell acquisition and data analysis were performed using Guava® and GuavaSoft (Merck) and then GraphPad Prism software.
Oxidative stress-dependent DNA damage analysis
To perform oxidative stress-dependent DNA damage analysis, cells from the MRC5 healthy cell line and U87-MG glioblastoma cell line were seeded in the same cell container, as describe above. Cells were irradiated at a target dose of , harvested one hour post-irradiation and immediately frozen in liquid nitrogen before being stored at . Oxidative stress-dependent DNA damage (8-hydroxyguanosine, 8-OHdG) measurement was performed using DNA damage ELISA kit (Enzo) following manufacturer protocol. Absorbance at was measured using an EnSpire Plate reader (PerkingElmer), and the data were analyzed using GraphPad Prism software.
Zebrafish holder
Owing to the limited available kinetic energy in the proton beam, a specific holder (Figure 15) for zebrafish embryos was designed and fabricated. The embryos are kept in a water reservoir before irradiation; the holder is kept horizontal (as in Figure 15-b) and the embryos are let sink in a deep notch on the top of the piston. At the moment of irradiation the piston is pushed upwards towards the Mylar water/air separation. In this condition embryos are confined in the inner cylinder, which makes possible to put the holder in vertical position (as in Figure 15-a) for installation on the irradiation beam-line.
Zebrafish embryos experiments
For in vivo studies, wild-type (WT) zebrafish were bred in LOB fish facility (École Polytechnique). All in vivo experiments on zebrafish were performed on embryos below days post-irradiation. Zebrafish embryos were obtained by natural spawning of WT fish. Fertilized WT zebrafish eggs were incubated at until post-irradiation. Irradiation was performed at a target dose of SP-LDP and FLASH (Oriatron, Institut Curie, as control) in zebrafish holder. For FLASH irradation, zebrafish holder was placed in a water tank. Embryos were fixed post-irradiation with a solution of paraformaldehyde ( final concentration for one hour) before microscopic analysis (Evos XL Core Cell Imaging System; Thermo Fisher Scientific). Fish length was measured using ImageJ 1.X. software.